# Mind the Error! Detection and Localization of Instruction Errors in Vision-and-Language Navigation

### v1

@@ Abstract
Vision-and-Language Navigation in Continuous Environments (VLN-CE) is one of the most intuitive yet challenging embodied AI tasks. Agents are tasked to navigate towards a target goal by executing a set of low-level actions, following a series of natural language instructions. All VLN-CE methods in the literature assume that language instructions are exact. However, in practice, instructions given by humans can contain errors when describing a spatial environment due to inaccurate memory or confusion. Current VLN-CE benchmarks do not address this scenario, making the state-of-the-art methods in VLN-CE fragile in the presence of erroneous instructions from human users. For the first time, we propose a novel benchmark dataset that introduces various types of instruction errors considering potential human causes. This benchmark provides valuable insight into the robustness of VLN systems in continuous environments. We observe a noticeable performance drop (up to $-25\%$) in Success Rate when evaluating the state-of-the-art VLN-CE methods on our benchmark. Moreover, we formally define the task of Instruction Error Detection and Localization, and establish an evaluation protocol on top of our benchmark dataset. We also propose an effective method, based on a cross-modal transformer architecture, that achieves the best performance in error detection and localization, compared to baselines. Surprisingly, our proposed method has revealed errors in the validation set of the two commonly used datasets for VLN-CE, i.e., R2R-CE and RxR-CE, demonstrating the utility of our technique in other tasks. Code and dataset will be made available upon acceptance at https://intelligolabs.github.io/R2RIE-CE
@@ I Introduction
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@
Figure 1: An agent navigates in a scene, following instructions expressed in natural language, for example “Exit the bathroom and go left (✓right), then turn left at the big clock and go into the bedroom and wait next to the bed.”. By just changing “right” to “left” in the instruction, the agent terminates the exploration in the wrong location, ignoring the fact that along the path it did not see the “big clock” (yellow arrow).
@@ II Related Works
Our work differs from the previous studies, as we investigate the instruction errors and their impact for navigation policies in the VLN task. We manipulate among a large set of tokens to inject instruction errors, instead of just masking certain tokens out as in [19]. We do not focus on path-ranking models [20], the domain gap towards unseen environments [20], and skill-specific interventions [21]. Instruction and Trajectory Alignment. Instruction and Trajectory Alignment is the ability to align semantically the concepts represented by these two modalities, an important aspect for obtaining a VLN policy. Huang et al. [22] introduces the Cross-Modal Alignment (CMA) task, i.e., discriminating instruction-trajectory path pairs from negative pairs, by maintaining the original instruction and altering the path sequence. Zhao et al. [23] perform a study on the quality of the instruction generated by VLN instruction generation models [13, 24] using human wayfinders. Moreover, they propose an Instruction-Trajectory compatibility model to classify high and low-quality instructions for trajectories from the R2R Val Unseen and Val Seen sets with the discrete VLN setting. These instructions are considered to be high quality if 2 out of 3 human wayfinders reached the goal. Differently from [22], they also consider instruction perturbations. In [25] the authors proposed a Contrastive Instruction-Trajectory Learning framework to boost the generalizability of the navigation policy. Their method selects sub-optimal trajectories by the hop distance from an anchor trajectory. Positive instructions are generated by substituting words with their synonym, inserting or substituting them with context or by back-translation; instead, negative ones are generated by splitting the original instruction into sub-instructions and shuffling or repeating randomly to become an intra-negative instruction. The authors assume that the augmented instructions should preserve the semantic information of the original ones.
@@ IV Benchmark
We establish the first benchmark dataset R2RIE-CE, short for R2R with Instruction Errors in Continuous Environments, for assessing VLN-CE methods with erroneous instructions and for the novel task of Detection and Localization of Instruction Errors. Our benchmark dataset is built on top of the R2R-CE [5] dataset, which serves for evaluating VLN-CE methods. The R2R-CE dataset consists of three splits, namely Train, Val Seen and Val Unseen. In particular, we will focus on the most challenging Val Unseen, which is composed of $1839$ episodes with new paths, instructions, and scenes that are not observed during training.
@@
In order to create erroneous instructions, we artificially inject various types of Instruction Errors (as defined in Sec. III) into the given natural language instructions in R2R-CE. Our error composition carefully considers error priors induced by human causes, including inaccurate scene memory and direction confusion. For each type of error, i.e., Direction, Object, Room, Room&Object and All, we create a corresponding validation set based on the Val Unseen validation split of R2R-CE. The original validation split can contain episodes that may not be appropriate for our evaluation. The natural language instruction should not be too short and the episodes should contain words relevant to the specific error setup. For example, episodes must have Direction words for the validation set with Direction Error, or episodes must contain Direction, Object, and Room words for the validation set with All Errors. Specifically, prior to constructing the validation set with each type of error, we first exclude episodes with a minimum length of less than $\tau$ words, and episodes lacking words relevant to the specific experiments, such as Direction error words for the Direction Error dataset. Following this filtering stage, we obtain a set of correct episodes $\mathcal{E_{C}}$ for each type of Instruction Error. Then, for each episode $e_{i}\in\mathcal{E_{C}}$, we create a corresponding erroneous episode in which we perturbed the instruction with the respective error. Specifically, (i) for the Direction Error, we consider the following set of directional words that occur frequently in the instructions: left/right, go down/go up, into/out of, forward/backward, inside/outside, go around/go back, leftmost/rightmost. We introduce the Direction Error by swapping the ground-truth direction with its antonym; (ii) For the Object Error, we first identify a set of common object categories $\mathcal{C}$ that frequently appear in a set of language instructions, excluding synonyms. Each object class $c_{i}\in\mathcal{C}$ is associated with a set of object classes $\mathcal{C}_{i}$ comprised of the classes that often co-locate in the same room. We introduce an Object Error by swapping the ground-truth object class $c_{i}$ to a random class $c_{j}\in\mathcal{C}_{i}$; (iii) For the Room Error, we consider the following set of rooms $\mathcal{R}$ that are common in indoor environments: kitchen, archway, bathroom, bedroom, gym, lounge, hallway, living room, office, dining room, laundry, restroom. Each room $r_{i}\in\mathcal{R}$ is associated with a set of rooms $\mathcal{R}_{i}$ that are often adjacent to $r_{i}$ [^2] The list is obtained by leveraging ConceptNet relations and manual curation. We introduce a Room Error by swapping the ground-truth room $r_{i}$ to a random room $r_{j}\in\mathcal{R}_{i}$; (iv) For the Room&Object Error, we introduce in an instruction both Room and Object Error; (v) For the All Error, we introduce in an instruction both Direction and Room&Object Error. The perturbed episodes will then be stored in the corresponding set of perturbed episodes, $\mathcal{E_{P}}$. For each perturbed episode $e_{i}\in\mathcal{E_{P}}$, both the error type and the position of the perturbed word are stored as metadata for the assessment of Detection and Localization of Instruction Errors. Eventually, for each type of instruction error, we obtain sets $\mathcal{E_{C}}$ and $\mathcal{E_{P}}$. Tab.I presents the statistics of the created validation set with each type of error.
@@ V Method
Figure 3: Architecture of our proposed IEDL model, representing the scenario depicted in Fig 1. The frozen policy $\pi$ follows Instruction $\Upsilon$, producing a sequence of observation $\mathcal{O}$. Then, a panoramic encoder and a language encoder produce, respectively, the trajectory visual features $\Gamma$ and instruction features $\Upsilon$. We then fed the trajectory set $\Gamma$ and $\Upsilon$ to a cross-modal multi-layer transformer to produce visual-language aligned features. Finally, two specialized heads perform Instruction Error Detection and Instruction Error Localization, respectively.
@@
Model structure. The full pipeline of IEDL is shown in Fig. 3. We define the instruction embeddings as $\Upsilon\in\mathbb{R}^{W\times D}$, namely the textual instruction $\mathcal{I}$ undergoes a tokenization and padding procedure, up to $W=80$ tokens [9, 10], with associated text BERT [26] embeddings. Moreover, we refer to $D$ as the dimensionality of the latent space. We consider the trajectory set of image embeddings $\Gamma=\{V_{1},...,V_{T}\}$ (where $V_{t}$ is the corresponding image embedding of visual observation $O_{t}$ defined in Section III) as the set containing the viewpoint panorama visual features $V_{t}\in\mathbb{R}^{D}$ (i.e., embeddings) of the node chosen at each time step $t$ by policy $\pi$, following instruction embeddings $\Upsilon$. As in [9], the embedding $V_{t}$ contains features extracted from a pre-trained Vision Transformer (ViT-B/16-CLIP [27]), which are then fed into a panoramic encoder [14] to obtain contextual view embeddings. For each episode $i$, we first collect the set of observation embeddings (i.e., trajectory) $\Gamma\in\mathbb{R}^{T\times D}$ from policy $\pi$. To retain positional information, we use sine and cosine positional encoding [28]. Similar to BERT [26] [CLS] token, we prepend a learnable [CLS] $\in\mathbb{R}^{D}$ embedding to set $\Gamma$, which will be later used for our downstream tasks. Then, trajectory set $\Gamma$ and instruction embeddings $\Upsilon$ are fused together by feeding them to a cross-modal multi-layer transformer with $k$ layers (see right part of Fig.3). We adopted an architecture similar to [29] but without the bi-directional cross-attention layer since our focus is not on having more than one cross-modal aligned output. For each layer of the transformer, we first perform cross-attention using trajectory set $\Gamma$ as a query ($Q$) and instruction embeddings $\Upsilon$ as a context ($KV$). Then, we perform self-attention followed by a feed-forward layer. The [CLS] token (enriched with visual-language fused features after the cross-modal transformer) is then fed to two classification heads that produce the output of our Instruction Error Detector & Localizer: (i) the trajectory-instruction matching head $f_{d}:\mathbb{R}^{D}\rightarrow\mathbb{R}$ (which implements function $d_{\pi}$ presented in Sec. III) predicts the alignment score $\sigma(a)\in[0,1]$, where $\sigma(\cdot)$ is the sigmoid operator; (ii) the error localization head $f_{l}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{W}$ (which implements function $l_{\pi}$ presented in Sec. III) predicts the token errors’ localization within the instruction. Both heads are implemented by a Multilayer Perceptron (MLP), a ReLU activation function, followed by a Layernorm layer and a final MLP.
@@ VI Experiments
(i) Random. Each instruction embeddings $\Upsilon_{i}$ for episode $i$ is randomly classified as correct or wrong. We then predict $J_{i}$ random token indices, of which each ^$\hat{\ell}^{i}_{j}\in[0,len(\Upsilon_{i})]$, where $J_{i}$ is the number of errors expected in the dataset (see Tab. I).
@@
Origin split Error type
Policy [9]
Random
CLIP Alignment
IEDL
SR normal-$\uparrow$ SPL normal-$\uparrow$ $\Delta_{SR}(\%)$ AUC normal-$\uparrow$ ATD normal-$\downarrow$ AUC normal-$\uparrow$ ATD normal-$\downarrow$ AUC normal-$\uparrow$ ATD normal-$\downarrow$
R2R-CE Val Unseen Direction 0.53 0.43 -18.64 0.50 10.54 0.50 11.05 0.59 7.74
${}^{\star}$ 0.58 0.49 -6.66 0.50 11.03 0.57 9.63 0.79 7.85
${}^{\star}$ 0.56 0.46 -8.47 0.51 10.94 0.59 8.76 0.74 10.14
Room&Object ${}^{\star}$ 0.57 0.47 -11.47 0.49 11.56 0.64 7.98 0.90 6.75
All ${}^{\star}$ 0.52 0.43 -30.64 0.51 12.22 0.63 8.68 0.93 5.68
Avg. 0.55 0.46 -15.17 0.50 11.26 0.59 9.22 0.82 7.46
@@
Does the instruction contain an error? In this test, the performance of the classification head $f_{d}$ of IEDL is evaluated against two baselines in terms of AUC scores. The results are reported in Tab. II. Here, the AUC scores of a random baseline are presented as a means to identify potential biases and establish a lower bound. Given that the ratio between $\mathcal{E_{C}}$ and $\mathcal{E_{P}}$ episodes is $50\%$, AUC for the random baseline is always $\sim 0.50$, as expected. The CLIP Alignment baseline is presented here as a simple method that reflects how humans would solve the same problem in real life: checking if the content of the instruction is grounded to the observations. Also, note that CLIP Alignment does not require training. CLIP Alignment seems to be effective in the presence of Object and Room types of errors, with a total $0.63$ of AUC score in Room&Object when both errors are present. Interestingly, CLIP Alignment performs on par with the random baseline for the Direction error type. The intuition is that, when errors are present near the end of the instruction, CLIP Alignment can still ground objects and room tokens to the observations (thus $\mathcal{K}\subseteq\mathcal{S}$), but may subsequently become lost due to the error. Finally, in Tab. II, we show our proposed IEDL method. We can see that IEDL achieves the best AUC in all the benchmarks by a large margin. A lower AUC of $0.59$ for IEDL also seems to highlight the challenges of the Direction error type. Finally, IEDL has a much higher average AUC score ($0.82$) compared to Random ($0.50$) and CLIP Alignment ($0.59$).
Can we localize the error? This experiment evaluates the ability of the localization head $f_{l}$ of IEDL to locate the errors within an instruction. The results in terms of ATD for the different baselines are also shown in Tab. II. The random baseline achieves a mean ATD of 11.26. The CLIP Alignment baseline performs better for all error types compared to random. The policy of considering as token index error $\hat{l}$ the instruction token that was not observed during navigation seems effective, especially in the Room&Object benchmark. Finally, the proposed IEDL achieves the best localization performance across all benchmarks, except for the Object, in which CLIP Alignment seems to be particularly effective. However, it should be noted that, in the Object case, the AUC score for CLIP Alignment is $0.59$ vs 0.74 of IEDL, which indicates the ability of CLIP Alignment in identifying only a subset of objects that have been swapped. Finally, it is worth noticing that the mean ATD of IEDL is $7.46$, which is close to the length of a typical sub-sentence within each instruction, showing that our method can be used to localize errors at the sub-sentence level.

### v2

@@ Abstract
Vision-and-Language Navigation in Continuous Environments (VLN-CE) is one of the most intuitive yet challenging embodied AI tasks. Agents are tasked to navigate towards a target goal by executing a set of low-level actions, following a series of natural language instructions. All VLN-CE methods in the literature assume that language instructions are exact. However, in practice, instructions given by humans can contain errors when describing a spatial environment due to inaccurate memory or confusion. Current VLN-CE benchmarks do not address this scenario, making the state-of-the-art methods in VLN-CE fragile in the presence of erroneous instructions from human users. For the first time, we propose a novel benchmark dataset that introduces various types of instruction errors considering potential human causes. This benchmark provides valuable insight into the robustness of VLN systems in continuous environments. We observe a noticeable performance drop (up to $-25\%$) in Success Rate when evaluating the state-of-the-art VLN-CE methods on our benchmark. Moreover, we formally define the task of Instruction Error Detection and Localization, and establish an evaluation protocol on top of our benchmark dataset. We also propose an effective method, based on a cross-modal transformer architecture, that achieves the best performance in error detection and localization, compared to baselines. Surprisingly, our proposed method has revealed errors in the validation set of the two commonly used datasets for VLN-CE, i.e., R2R-CE and RxR-CE, demonstrating the utility of our technique in other tasks.
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Figure 1: An agent navigates in a scene, following instructions expressed in natural language, for example “Exit the bathroom and go left (✓right), then turn left at the big clock and go into the bedroom and wait next to the bed.” By just changing “right” to “left” in the instruction, the agent terminates the exploration in the wrong location, ignoring the fact that along the path it did not see the “big clock” (yellow arrow).
@@ II Related Works
Our work differs from the previous studies, as we investigate the instruction errors and their impact for navigation policies in the VLN task. We manipulate among a large set of tokens to inject instruction errors, instead of just masking certain tokens out as in [19]. We do not focus on path-ranking models [20], the domain gap towards unseen environments [20], and skill-specific interventions [21].
Instruction and Trajectory Alignment. Instruction and Trajectory Alignment is the ability to align semantically the concepts represented by these two modalities, an important aspect for obtaining a VLN policy. Huang et al. [22] introduces the Cross-Modal Alignment (CMA) task, i.e., discriminating instruction-trajectory path pairs from negative pairs, by maintaining the original instruction and altering the path sequence. Zhao et al. [23] perform a study on the quality of the instruction generated by VLN instruction generation models [13, 24] using human wayfinders. Moreover, they propose an Instruction-Trajectory compatibility model to classify high and low-quality instructions for trajectories from the R2R Val Unseen and Val Seen sets with the discrete VLN setting. These instructions are considered to be high quality if 2 out of 3 human wayfinders reached the goal. Differently from [22], they also consider instruction perturbations. In [25] the authors proposed a Contrastive Instruction-Trajectory Learning framework to boost the generalizability of the navigation policy. Their method selects sub-optimal trajectories by the hop distance from an anchor trajectory. Positive instructions are generated by substituting words with their synonym, inserting or substituting them with context or by back-translation; instead, negative ones are generated by splitting the original instruction into sub-instructions and shuffling or repeating randomly to become an intra-negative instruction. The authors assume that the augmented instructions should preserve the semantic information of the original ones.
@@ IV Benchmark
We establish the first benchmark dataset R2RIE-CE, short for R2R with Instruction Errors in Continuous Environments, for assessing VLN-CE methods with erroneous instructions and for the novel task of Detection and Localization of Instruction Errors. Our benchmark dataset is built on top of the R2R-CE [5] dataset, which serves for evaluating VLN-CE methods. The R2R-CE dataset is split into Train, Val Seen and Val Unseen. We will use Train set for training data and focus on the most challenging Val Unseen for validation, which has $1839$ episodes with new paths, instructions, and scenes that are not observed during training.
@@
In order to create erroneous instructions, we artificially inject various types of Instruction Errors (as defined in Sec. III) into the given natural language instructions in R2R-CE. Our error composition carefully considers error priors induced by human causes, including inaccurate scene memory and direction confusion. For each type of error, i.e., Direction, Object, Room, Room&Object and All, we create a corresponding validation set based on the Val Unseen validation split of R2R-CE. The original validation split can contain episodes that may not be appropriate for our evaluation. The natural language instruction should not be too short and the episodes should contain words relevant to the specific error setup. For example, episodes must have Direction words for the validation set with Direction Error, or episodes must contain Direction, Object, and Room words for the validation set with All Errors. Specifically, prior to constructing the validation set with each type of error, we first exclude episodes with a minimum length of less than $\tau$ words, and episodes lacking words relevant to the specific experiments, such as Direction error words for the Direction Error dataset. Following this filtering stage, we obtain a set of correct episodes $\mathcal{E_{C}}$ for each type of Instruction Error.
Then, for each episode $e_{i}\in\mathcal{E_{C}}$, we create a corresponding erroneous episode in which we perturbed the instruction with the respective error. Specifically, (i) for the Direction Error, we consider the following set of directional words that occur frequently in the instructions: left/right, go down/go up, into/out of, forward/backward, inside/outside, go around/go back, leftmost/rightmost. We introduce the Direction Error by swapping the ground-truth direction with its antonym; (ii) For the Object Error, we first identify a set of common object categories $\mathcal{C}$ that frequently appear in a set of language instructions, excluding synonyms. Each object class $c_{i}\in\mathcal{C}$ is associated with a set of object classes $\mathcal{C}_{i}$ comprised of the classes that often co-locate in the same room. We introduce an Object Error by swapping the ground-truth object class $c_{i}$ to a random class $c_{j}\in\mathcal{C}_{i}$; (iii) For the Room Error, we consider the following set of rooms $\mathcal{R}$ that are common in indoor environments: kitchen, archway, bathroom, bedroom, gym, lounge, hallway, living room, office, dining room, laundry, restroom. Each room $r_{i}\in\mathcal{R}$ is associated with a set of rooms $\mathcal{R}_{i}$ that are often adjacent to $r_{i}$ [^2] The list is obtained by leveraging ConceptNet relations and manual curation. We introduce a Room Error by swapping the ground-truth room $r_{i}$ to a random room $r_{j}\in\mathcal{R}_{i}$; (iv) For the Room&Object Error, we introduce in an instruction both Room and Object Error; (v) For the All Error, we introduce in an instruction both Direction and Room&Object Error. The perturbed episodes will then be stored in the corresponding set of perturbed episodes, $\mathcal{E_{P}}$. For each perturbed episode $e_{i}\in\mathcal{E_{P}}$, both the error type and the position of the perturbed word are stored as metadata for the assessment of Detection and Localization of Instruction Errors. Eventually, for each type of instruction error, we obtain sets $\mathcal{E_{C}}$ and $\mathcal{E_{P}}$. Tab.LABEL:tab:statistic presents the statistics of the created validation set with each type of error.
@@ V Method
Figure 3: Architecture of our proposed IEDL model, representing the scenario depicted in Fig 1. The frozen policy $\pi$ follows Instruction $\Upsilon$, producing a sequence of observation $\mathcal{O}$. Then, a panoramic encoder and a language encoder produce, respectively, the trajectory visual features $\Gamma$ and instruction features $\Upsilon$. We then feed the trajectory set $\Gamma$ and $\Upsilon$ to a cross-modal multi-layer transformer to produce visual-language aligned features. Finally, two specialized heads perform Instruction Error Detection and Instruction Error Localization, respectively.
@@
Model structure. The full pipeline of IEDL is shown in Fig. 3. We define the instruction embeddings as $\Upsilon\in\mathbb{R}^{W\times D}$, namely the textual instruction $\mathcal{I}$ undergoes a tokenization and padding procedure, up to $W=80$ tokens [9, 10], with associated text BERT [26] embeddings. Moreover, we refer to $D$ as the dimensionality of the latent space. We consider the trajectory set of image embeddings $\Gamma=\{V_{1},...,V_{T}\}$ (where $V_{t}$ is the corresponding image embedding of visual observation $O_{t}$ defined in Section III) as the set containing the viewpoint panorama visual features $V_{t}\in\mathbb{R}^{D}$ (i.e., embeddings) of the node chosen at each time step $t$ by policy $\pi$, following instruction embeddings $\Upsilon$. As in [9], the embedding $V_{t}$ contains features extracted from a pre-trained Vision Transformer (ViT-B/16-CLIP [27]), which are then fed into a panoramic encoder [14] to obtain contextual view embeddings. For each episode $i$, we first collect the set of observation embeddings (i.e., trajectory) $\Gamma\in\mathbb{R}^{T\times D}$ from policy $\pi$. To retain positional information, we use sine and cosine positional encoding. Similar to BERT [26] [CLS] token, we prepend a learnable [CLS] $\in\mathbb{R}^{D}$ embedding to set $\Gamma$, which will be later used for our downstream tasks. Then, trajectory set $\Gamma$ and instruction embeddings $\Upsilon$ are fused together by feeding them to a cross-modal multi-layer transformer with $k$ layers (see right part of Fig.3). We adopted an architecture similar to [28] but without the bi-directional cross-attention layer since our focus is not on having more than one cross-modal aligned output. For each layer of the transformer, we first perform cross-attention using trajectory set $\Gamma$ as a query ($Q$) and instruction embeddings $\Upsilon$ as a context ($KV$). Then, we perform self-attention followed by a feed-forward layer. The [CLS] token (enriched with visual-language fused features after the cross-modal transformer) is then fed to two classification heads that produce the output of our Instruction Error Detector & Localizer: (i) the trajectory-instruction matching head $f_{d}:\mathbb{R}^{D}\rightarrow\mathbb{R}$ (which implements function $d_{\pi}$ presented in Sec. III) predicts the alignment score $\sigma(a)\in[0,1]$, where $\sigma(\cdot)$ is the sigmoid operator; (ii) the error localization head $f_{l}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{W}$ (which implements function $l_{\pi}$ presented in Sec. III) predicts the token errors’ localization within the instruction. Both heads are implemented by a Multilayer Perceptron (MLP), a ReLU activation function, followed by a Layernorm layer and a final MLP.
@@ VI Experiments
(i) Random. Each instruction embeddings $\Upsilon_{i}$ for episode $i$ is randomly classified as correct or wrong. We then predict $J_{i}$ random token indices, of which each ^$\hat{\ell}^{i}_{j}\in[0,len(\Upsilon_{i})]$, where $J_{i}$ is the number of errors expected in the dataset (see Tab. LABEL:tab:statistic).
@@
Origin split Error type Policy [9] Random CLIP Alignment IEDL
SR $\uparrow$ SPL $\uparrow$ $\Delta_{SR}(\%)$ AUC $\uparrow$ ATD $\downarrow$ AUC $\uparrow$ ATD $\downarrow$ AUC $\uparrow$ ATD $\downarrow$
R2R-CE Val Unseen Direction 0.53 0.43 -18.64 0.50 10.54 0.50 11.05 0.58 8.13
Room 0.58 0.49 -6.66 0.50 11.03 0.57 9.63 0.80 7.73
Object 0.56 0.46 -8.47 0.51 10.94 0.59 8.76 0.74 9.21
Room&Object 0.57 0.47 -11.47 0.49 11.56 0.64 7.98 0.91 7.34
All 0.52 0.43 -30.64 0.51 12.22 0.63 8.68 0.94 6.14
Avg. 0.55 0.46 -15.17 0.50 11.26 0.59 9.22 0.79 7.71
@@
Does the instruction contain an error? In this test, the performance of the classification head $f_{d}$ of IEDL is evaluated against two baselines in terms of AUC scores. The results are reported in Tab. II. Here, the AUC scores of a random baseline are presented as a means to identify potential biases and establish a lower bound. Given that the ratio between $\mathcal{E_{C}}$ and $\mathcal{E_{P}}$ episodes is $50\%$, AUC for the random baseline is always $\sim 0.50$, as expected. The CLIP Alignment baseline is presented here as a simple method that reflects how humans would solve the same problem in real life: checking if the content of the instruction is grounded to the observations. Also, note that CLIP Alignment does not require training. CLIP Alignment seems to be effective in the presence of Object and Room types of errors, with a total $0.64$ of AUC score in Room&Object when both errors are present. Interestingly, CLIP Alignment performs on par with the random baseline for the Direction error type. The intuition is that, when errors are present near the end of the instruction, CLIP Alignment can still ground objects and room tokens to the observations (thus $\mathcal{K}\subseteq\mathcal{S}$), but may subsequently become lost due to the error. Finally, in Tab. II, we show our proposed IEDL method. We can see that IEDL achieves the best AUC in all the benchmarks by a large margin. A lower AUC of $0.58$ for IEDL also seems to highlight the challenges of the Direction error type. Finally, IEDL has a much higher average AUC score ($0.79$) compared to Random ($0.50$) and CLIP Alignment ($0.59$).
Can we localize the error? This experiment evaluates the ability of the localization head $f_{l}$ of IEDL to locate the errors within an instruction. The results in terms of ATD for the different baselines are also shown in Tab. II. The random baseline achieves a mean ATD of 11.26. The CLIP Alignment baseline performs better for all error types compared to random. The policy of considering as token index error $\hat{l}$ the instruction token that was not observed during navigation seems effective, especially in the Room&Object benchmark. Finally, the proposed IEDL achieves the best localization performance across all benchmarks, except for the Object, in which CLIP Alignment seems to be particularly effective. However, it should be noted that, in the Object case, the AUC score for CLIP Alignment is $0.59$ vs 0.74 of IEDL, which indicates the ability of CLIP Alignment in identifying only a subset of objects that have been swapped. Finally, it is worth noticing that the mean ATD of IEDL is $7.71$, which is close to the length of a typical sub-sentence within each instruction, showing that our method can be used to localize errors at the sub-sentence level.